A helium nucleus is sent at high speed toward a metal sphere with a charge of 1.5 nC what is the force acting on the helium nucleus when it is 3.5 mm away from the metal sphere


1
Expert's answer
2021-02-18T18:53:33-0500

Answer

Given data

Q=ze=2e charge of helium atom

Q'=1.5 nC

r=3.5 mm

the force acting on the helium nucleus 

"F=\frac{KQQ'}{r^2}"

"=\frac{9\times10^9\times2\times1.6\times10^{-19}\times1.5\times10^{-9}}{3.5\times3.5\times \times10^{-6} }"

"=3.5\times10^{-13}N"